Ralph Chase plans to sell a piece of property for ​$145000. He wants the money to be paid off in two waysminus−a ​short-term note at 12​% interest and a​ long-term note at 9% interest. Find the amount of each note if the total annual interest paid is ​$15450
he amount of the 12​% note is ​$____ and the amount of the 9​% note is ​$___

Answer by Theo(13342)

the short term note pays 12% interest per year.
the long term note pays 9% interest per year.

the total interest per year is 145,000.

let x = the amount of the short term loan.
let y = the amount of the long term loan.

x + y = 145000

.12x + .09y = 15450

you have 2 equations that needs to be solved simultaneously.

they are:

x + y = 145000
.12x + .09y = 15450


assuming you know how to solve equations simultaneously, i'll just do this graphically and show you the result.

the intersections of the lines of both equations shows the solution.

that occurs when x = 80000 and y = 65000

80000 + 65000 = 145000

.12*80000 + .09*65000 = 15450

this is what the graph looks like:

$$$

if you do not know how to solve equations simultaneously, here's a good reference.

http://www.purplemath.com/modules/systlin1.htm